Teal Inc. loans money to John Kruk Corporation in the amount of $1,008,000. Teal accepts an 8% note due in 7 years with interest payable semiannually. After 2 years (and receipt of interest for 2 years), Teal needs money and therefore sells the note to Chicago National Bank, which demands interest on the note of 10% compounded semiannually. What is the amount Teal will receive on the sale of the note? (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)
Answer #1
Amount received on the sale of the note 930161
Workings:
Semi-annual interest 40320 =1008000*8%*6/12
X PV factor 7.72173 =(1-(1.05)^-10)/0.05
Present value of interest 311340
Principal 1008000
X PV factor 0.61391 =1/1.05^10
Present value of principal 618821
Amount received on the sale of the note 930161 =311340+618821